OpenStudy (anonymous):

Consider the functions f(x) = 3x2 and g(x) = 2(x + 9). Find f[g(−7)].

OpenStudy (anonymous):

well, first get the (-7) value of g by plugging -7 into the equation for g:\[g(-7) = 2\cdot((-7) + 9) = 2\cdot 2 = 4\]Now we know, as \[g(-7) = 4 \implies f(g(-7)) = f(4)\] So we plug x = 4 into f and are done \[f(4) = 3\cdot 4^2 = 3\cdot 16 = 48\]